OpenStudy (anonymous):

to solve the equation f(x)=0 by Newton's Method we start with a good initial guess x0 and then run the iteration xn+1=xn- [f(xn)/f'(xn)], n = 0, 1, 2,... until we get an approximation xn+1. Suppose that you want to compute the cube root of 4 by solving the equation x^3-4=0. Since 1^3=1 and 2^3=8, let's start with x0=1.5 Then x1=?, x2=?, x3=? (with at least 6 correct digits beyond the decimal point.)

OpenStudy (anonymous):

\[f(x)=x^3-4\] \[f'(x)=3x^2\] \[x_{n+1}=x_{n}-\frac{f(x_n)}{f'(x_n)}\] \[x_{n+1}=x_n-\frac{x_n^3-4}{3x_n^2}\]

OpenStudy (anonymous):

we can either do the algebra first and then plug in the numbers or just plug them in, it doesn't matter.

OpenStudy (anonymous):

by "algebra" i mean write \[x_{n+x}=\frac{3x_n^3-x_n^3+4}{3x_n^3}=\frac{2x_n^3+4}{3x_n^2}\]

OpenStudy (anonymous):

that should be \[x_{n+1}\] not \[x_{n+x}\]

OpenStudy (anonymous):

now plug in 1.5 and see what you get. then do it again with the result. a spread sheet would be nice, but a calculator will do.

OpenStudy (anonymous):

Got it. Thanks, sat :]